# Taming systematics in distance and inclination measurements with gravitational waves: role of the detector network and higher-order modes

### v1

@@ I Introduction
An accurate measurement of these parameters is crucial in applications ranging from cosmology to nuclear physics. For example, a combination of the luminosity distance measurement from a GW signal and the redshift measurement from an electromagnetic counterpart, from the same event, can be used to infer the value of the Hubble constant [28, 29], as done for GW170817 [30]. More in general, future GW observations can play a crucial role in testing the current $\Lambda$CDM cosmological model and in searching for new physics, because of their powerful property of being directly sensitive to the luminosity distance. This implies the possibility of measuring the Hubble constant with a statistical approach even when an electromagnetic counterpart is not present [31, 32, 33]. Distance measurements are also pivotal in constraining the equation of state of neutron stars [34, 35]. Indeed, while GW detections of BNSs allow for measurements of the neutron star tidal deformability [36, 37, 38] and the masses in the detector frame, in order to recover the equation of state one needs information about the masses in the source frame, which inevitably requires knowledge of the source redshift. The latter cannot be directly extracted from the waveform. However, if an accurate measurement of the distance is available, the redshift can be inferred assuming a cosmological model [39] or, for low-redshift sources, simply the Hubble law.
@@
All these examples highlight the utmost importance of measuring the luminosity distance $d_{L}$ and the inclination angle $\iota$ as accurately and precisely as possible with GWs. However, a degeneracy between $d_{L}$ and $\iota$ [48] makes placing constraints over the two parameters difficult. The quasi indistinguishability of the polarization amplitudes in a broad region for $\iota\gtrsim 0\degree$ is such that low-inclination systems are mistaken for closer, more inclined ones since the distance acts as a scale factor in the waveform. The issue is exacerbated by the selection effect for which face-on (face-off) binaries, namely ones with $\iota=0\degree\,(180\degree)$, are more likely to be detected [49]. Additionally, the inability of current instruments to detect the two polarizations with the same accuracy [48] might extend the problem to higher inclination angles [50].
In principle, there are various ways to break this degeneracy: by using the measurements of $z$$d_{L}$ and inclination from electromagnetic counterparts, as for GW170817 [51, 52], by detecting higher-order modes [53, 54], by measuring precession effects [55, 56] or, for BNSs, through the binary Love relations between the progenitor tidal deformabilities [57].
In this work, we present a comprehensive exploration of the effects of the distance-inclination degeneracy on GW measurements using synthetic data for a network of current and planned GW detectors, specifically for ET. We perform Bayesian inference using the public software Bilby [58]. We quantify the role of the network alignment factor, detector sensitivity, and waveform higher-order modes in mitigating or breaking the $d_{L}$$\iota$ degeneracy, which is almost exact in the low-inclination limit.333 Reference [50] performed a similar analysis for what concerns the alignment factor of the network. In addition to extending their analysis, here we also also investigate the role of the high-order modes and the inclination angles.
@@ III.1 Origin of the degeneracy
The two polarizations may be written, at Newtonian level (i.e. 00-post-Newtonian) as:
@@ III.2.1 Alignment factor
$\begin{split}\mathcal{F}^{\,\text{dom}}[f]&=\mathcal{U}[f]\mathcal{F}[f]\mathcal{U}^{T}[f]=\begin{pmatrix}g_{++}[f]&0\\ 0&g_{\times\times}[f]\end{pmatrix}\\ &=g_{++}[f]\begin{pmatrix}1&0\\ 0&\alpha^{2}[f]\end{pmatrix},\end{split}$ (18)
@@
and where the subscripts $(1,2)$ represent the two polarizations (+,×)(+$\times$)( + , × ) in the new frame,
@@ III.2.2 High-order modes: the multipolar structure of gravitational radiation
$\begin{split}h_{+}^{0.5}(t)=&\frac{G\nu}{4d_{L}c^{2}}\,\delta m\,x^{3/2}\sin\iota\>\,\times\\ &\times\,\left[(\cos^{2}\iota+5)\cos\phi-9(1+\cos^{2}\iota)\cos 3\phi\right]\\ h_{\times}^{0.5}(t)=&\frac{3G\nu}{4d_{L}c^{2}}\,\delta m\,x^{3/2}\sin 2\iota\,\left[\sin\phi-3\sin 3\phi\right]\,,\end{split}$ (25)
@@
IV Results for injection simulations
@@ IV.1 Setup
In this section, we investigate how to constrain or break the $d_{L}$$\iota$ degeneracy, with Bayesian inference simulations performed with Bilby [58]. The baseline for our results is the set of simulations performed in Ref. [48], based on 4-parameter CBC Bayesian inference. The authors sampled over coalescence phase, polarization angle, luminosity distance and inclination, keeping fixed the coalescence time, the sky location, and the intrinsic parameters (with spins set to zero). The first generalization we bring is the inclusion of the mass intrinsic parameters among the sampled ones, albeit still maintaining the zero-spin configuration and the fixed sky location. In all the results that we present, we therefore neglect the effect of precession in the waveform, by imposing aligned spins injected to zero; we address the effect of precessing spins in the degeneracy in a future work [71]. Further, sampling over the sky-location parameters would indeed lead to larger uncertainties in distance and inclination measurements arising from varying detector sensitivities over the sky, while the focus of our investigating is the accuracy of measurements of distance and inclination in relation to the intrinsic degeneracy between the two quantities. Moreover, whenever a compact binary source with at least one neutron star is being considered, an electromagnetic counterpart observation could help in accurately localizing the source (possibly also constraining the inclination angle) thus reconducting the uncertainty problem back to the one of our interest [72].
@@
[…]
which is the probability distribution of Eq. (2) marginalized over all parameters except luminosity distance and inclination. This can be further marginalize to produce two single-parameter distributions. The quality of recovering each parameter is encoded in the accuracy that can be represented by a certain “distance” between the median of the 1-D marginalized distribution and the injected parameter, that is, the one characterizing the simulated source. However, this distance should provide statistical information, since we want to know if the true parameter is contained within a certain credible interval. This also depends on the measurement precision, which is reflected in the spread of the distribution. Given a parameter $\theta^{i}$, a good statistical distance can be defined as the absolute difference between the quantile order related to the injected value and $0.5$, that is, the quantile order of the median value. Given a simulation with injected parameter $\theta^{i}_{\text{inj}}$ and estimated median $\theta^{i}_{\text{med}}$, we thus define the following statistical distance:
@@
where $p(\theta^{i}|s)$ is the posterior marginalized over all the parameters except $\theta_{i}$, and $P(\theta^{i}|s)$ is the cumulative of $p(\theta^{i}|s)$. This quantity approaches zero aas the accuracy of the estimate increases. Conversely, bad recoveries would push this quantity toward $0.5$. Whenever an injected parameter is at the boundary of the typical prior support for that variable, it does not make sense to use this distance, since it will always return $0.5$. For instance, when injecting with an inclination angle of $0\degree$, a more indicative quantity might be the, e.g., the 90% credible upper limit. [^7] Or the lower limit, depending on the cases.
@@ IV.2 Role of the alignment factor
$\mathcal{M}$ Uniform in $[0.95\mathcal{M}_{\text{inj}},1.05\mathcal{M}_{\text{inj}}]$
$q$ Uniform in [0.5, 1]0.51[0.5,\,1][ 0.5 , 1 ]
@@
We choose slightly different component masses for computational reasons, and we sample over (detector-frame) chirp mass and mass ratio, as these are estimated more directly from GW data. Chirp mass, in particular, is the best-estimated parameter and therefore, to make the simulations faster, we choose a uniform prior centered in its injected value and with a 10% width around it. The injected mass ratio is close to unity, and the prior is taken to be uniform between $0.5$ and $1$. We provide priors for the (dimensionless) spin components aligned with the orbital angular momentum according to Ref. [73]. The extrinsic parameters listed in Tab. 1 are unchanged throughout these simulations.
The injected inclination angle $\iota$ varies over the simulations: it is taken to give face-on, edge-on, or intermediate-inclination binaries. The prior associated to this quantity is uniform in the cosine of $\iota$. Finally, we use a uniform-in-volume prior for the luminosity distance, which is degenerate with the inclination angle; the support for this prior abundantly contains the injected distance, but it necessarily varies with the injected $d_{L}$ as this is modified to yield different values of optimal SNR. Throughout this work, we perform zero-noise injections.
@@ IV.2.1 Two Advanced LIGO detectors
We now compare these results to the ones for $\hat{\rho}=40$ (bottom row in Fig. 4). One can immediately check that the poor recovery for the two systems at $\hat{\rho}=12$ had different causes. The face-on binary on the left maintains a degenerate profile which is smoother this time around, with a small enhancement in the precision [^8] We can quantify the precision of the estimate as the ratio between $\Delta d_{L}$, that is the larger part of the 90% credible interval, given the asymmetry of the distributions, and the median value $d_{L}^{\text{med}}$. We call it $\Delta d_{L}/d_{L}$ for simplicity. $\Delta d_{L}/d_{L}$ from $36\%$ to $26\%$, and with the upper limit on the inclination angle dropping from $70\degree$ to $55\degree$. The right plot shows the edge-on case, and proves that $\hat{\rho}=40$ is enough to properly appreciate the difference between the two polarizations and to improve the recovery: the injected inclination is almost coincident with the median value, and is included in a 90% credible interval $[84\degree,96\degree]$. As far as the distance is concerned, we find $\delta_{\text{st}}(d_{L})=0.26$, so that the injection is inside the 68% credible interval, with a moderate precision of $\Delta d_{L}/d_{L}=16\%$, which would further improve with a larger SNR.
@@ IV.2.3 The importance of signal-to-noise ratio and alignment factor
The other key factor in breaking the degeneracy is represented by the alignment factor $\alpha$, the impact of which has been widely discussed. We perform three simulations of the same system, with and inclination of 66°and detected at an optimal SNR $\hat{\rho}=18.5$ by three different networks: we add ET to the two networks already used so far, but at a different location with respect the one shown in Tab. 1, chosen in such a way to obtain a frequency-averaged $\alpha\simeq 0.9$. This allows us to have a complete range from $0.1$ to $0.9$. Figure 7 reports the contours of the 68% and 90% credible regions. The injection is inside the 90% credible region in the low-$\bar{\alpha}$ case and enters the 68% region when $\bar{\alpha}$ increases. The main differences between the three configurations can be observed in Tab. 3. The precision of the estimates, shown in the third and sixth columns, improves for the inclination and seems to worsen for the distance, but this is not an issue, since we worked at fixed SNR. [^9] This implies that the values of the distance were different such to have the same SNR. On the other hand, the accuracy, described by the other columns, improves sensibly, with the difference between the injected inclination and its median estimated value that drops from 48% to 11%. The values are slightly higher for the distance as discussed previously, but the enhancements are significant: from 65% with Advanced LIGO HL to 26% with ET. We have thus proved that the configurations in which the inclination angle is quite far from zero are manageable with a combination of $\bar{\alpha}$ and $\hat{\rho}$ achievable by the upcoming observing runs. However, the recovery of less inclined systems remains a problem difficult to solve even for 3G detectors, unless the SNRs reach very high and rare values [50], or minor effects contribute to distinguishing the two polarization amplitudes and breaking the correlation.
@@ IV.3.2 Low-mass-ratio binaries: an optimistic configuration
We now analyze the case of $\hat{\rho}=40$, when the SNR is twice the one seen so far, using IMRPhenomXHM for the recovery. The 2D marginalized posteriors for $d_{L}$ and $\iota$ are shown in Fig. 9. As expected, the credible regions shrink by a factor $\sim(40/20)^{2}=4$. The difference between injected and recovered distance is about $0.4\%$ for the higher SNR scenario, while the upper limit on the inclination angle drops to $\sim 10\degree$ in this same case. The statistical distance $\delta_{\text{st}}(d_{L})$ remains almost unchanged, while the precision is more than doubled. We summarize the results of these simulations in Tab. 4.
@@ IV.3.3 The role of ET design for low-mass-ratio binaries
    a 2L configuration with 15 km arms, with a relative orientation of 45°$\degree$, with the same locations of the previous configuration.
@@
Figure 10: Joint $d_{L}$$\iota$ posterior distributions for different ET designs, with injected $\iota=0\degree$ (left panels) and $\iota=66\degree$ (right panels); the recovery in the top and bottom panels is performed without and with higher-order modes, respectively. The black lines mark the injected values. The injected luminosity distance is fixed to 4Gpc, while the SNR values are reported in Tab. 5.
For each design, we set the inclination angle of the simulated source to two different values: $\iota\in\{0\degree,66\degree\}$. In both cases, we perform a recovery with and without high-order modes and the distance and sky location are kept unchanged. The corresponding optimal SNR values are reported in Tab. 5.
@@ IV.3.4 Next-to-unity mass ratio binaries I: a more challenging configuration
We begin by taking face-on BBHs with injected (detector-frame) masses of $36\,M_{\odot}$ and $29\,M_{\odot}$, close to the GW150914 values; we also perform a simulation at the same mass ratio but with neutron-star-compatible masses in order to confirm the similarity of the outcomes. For the BBHs, we varied the SNR from 20 to 150. As expected, the impact of high-order modes on the recovery is reduced, given that they are suppressed in the injected signal by the quasi-symmetry of the masses and therefore higher SNRs are appreciate them. The trend of the 90% credible upper limit on the inclination angle as a function of the optimal SNR is reported in the upper panel of Fig. 11, while the lower panel is the posterior distribution for the specific simulation with an optimal SNR of 100. The upper limit value on the inclination angle for $\hat{\rho}=150$ is almost the same as the one estimated at an SNR of 20 but with $q_{\rm inj}=0.1$ (see Tab. 4). The percentage difference between injected distance and recovered median value passes from 17% to 3%. However, the posterior distribution over the distance, with increasing SNR, does not tend to a normal distribution, but it shrinks asymmetrically toward the injected value that remains out of the 68% credible interval, as shown in the lower panel of Fig. 11. For this reason, the statistical distance $\delta_{\text{st}}(d_{L})$ does not improve.
Figure 12 reports the difference induced by the usage of IMRPhenomXHM or IMRPhenomXP for the recovery, to emphasize that higher-order modes remain important, but their impact is limited by the high mass ratio. The 90% contours are related to an SNR $\hat{\rho}=100$, and the key differences are summarized in Tab. 6. This table shows that considering high-order modes remains beneficial for both accuracy and precision, but the comparison with Tab. 4 highlights that a lower mass ratio is crucial.
Moving to BNS systems, we considered masses of $1.4\,M_{\odot}$ and $1.1\,M_{\odot}$, maintaining the same optimal SNR $\hat{\rho}=100$, that for ET corresponds to an injected luminosity distance of $\sim 200$ Mpc. We again used the two different approximants for the recoveries in two separate runs to ensure consistency with the previous test. We do not report the marginalized posteriors, which are analogous to the ones in Fig. 11, but we rather present Tab. 6 to be compared to Tab. 7.

### v2

# iano Frattale Mascioli0000-0002-0155-3833

@@ I Introduction
[…]
@@
All these examples highlight the utmost importance of measuring the luminosity distance $d_{L}$ and the inclination angle $\iota$ as accurately and precisely as possible with GWs. However, a degeneracy between $d_{L}$ and $\iota$ [48] makes placing constraints over the two parameters difficult. The quasi indistinguishability of the polarization amplitudes in a broad region for $\iota\gtrsim 0\degree$ is such that low-inclination systems are mistaken for closer, more inclined ones since the distance acts as a scale factor in the waveform. The issue is exacerbated by the selection effect for which face-on (face-off) binaries, namely ones with $\iota=0\degree\,(180\degree)$, are more likely to be detected [49]. Additionally, the inability of current instruments to detect the two polarizations with the same effectiveness [48] might extend the problem to higher inclination angles [50].
[…]
In this work, we present a comprehensive exploration of the effects of the distance-inclination degeneracy on GW measurements using synthetic data in the zero-noise configuration, for a network of current and planned GW detectors, specifically for ET. We perform Bayesian inference using the public software Bilby [58]. We quantify the role of the network alignment factor, detector sensitivity, and waveform higher-order modes in mitigating or breaking the $d_{L}$$\iota$ degeneracy, which is almost exact in the low-inclination limit.333 Reference [50] performed a similar analysis for what concerns the alignment factor of the network. In addition to extending their analysis, here we also also investigate the role of the high-order modes and the inclination angles.
@@ III.1 Origin of the degeneracy
The two polarizations may be written, at Newtonian level (i.e. 0-post-Newtonian) as:
@@ III.2.1 Alignment factor
[…]
@@
and where the subscripts $(1,2)$ represent the two polarizations $(+,\times)$ in the new frame,
@@ III.2.2 High-order modes: the multipolar structure of gravitational radiation
[…]
@@
IV Results for zero-noise injection simulations
@@ IV.1 Setup
In this section, we investigate strategies to break or mitigate the $d_{L}$$\iota$ degeneracy by performing a zero-noise injection-recovery study using Bayesian inference simulations carried out with Bilby [58]. The baseline for our results is the set of simulations performed in Ref. [48], based on 4-parameter CBC Bayesian inference. The authors sampled over coalescence phase, polarization angle, luminosity distance and inclination, keeping fixed the coalescence time, the sky location, and the intrinsic parameters (with spins set to zero). The first generalization we bring is the inclusion of the mass intrinsic parameters among the sampled ones, albeit still maintaining the zero-spin configuration and the fixed sky location. In all the results that we present, we therefore neglect the effect of precession in the waveform, by imposing aligned spins injected to zero; we address the effect of precessing spins in the degeneracy in a future work [71]. Further, sampling over the sky-location parameters would indeed lead to larger uncertainties in distance and inclination measurements arising from varying detector sensitivities over the sky, while the focus of our investigating is the accuracy of measurements of distance and inclination in relation to the intrinsic degeneracy between the two quantities. Moreover, whenever a compact binary source with at least one neutron star is being considered, an electromagnetic counterpart observation could help in accurately localizing the source (possibly also constraining the inclination angle) thus reconducting the uncertainty problem back to the one of our interest [72].
@@
$p(d_{L},\iota|s)=\int p(d_{L},\iota,\bm{\xi}|s)\,d\bm{\xi}\,$, (29)
which is the probability distribution of Eq. (2) marginalized over all parameters except luminosity distance and inclination. This can be further marginalized to produce two single-parameter distributions. The quality of recovering each parameter is encoded in the accuracy that can be represented by a certain “distance” between the median of the 1-D marginalized distribution and the injected parameter, that is, the one characterizing the simulated source. Given the use of injections in zero-noise realizations, the accuracy of parameter recovery is influenced by the choice of prior distributions and by the marginalization over the remaining parameters [^7] In a zero-noise injection-recovery simulation using the same waveform model for both injection and recovery — as is always the case in our setup — the injected parameters correspond to the maximum of the full $N$-dimensional likelihood. However, this is not necessarily true for the marginalized likelihoods. Moreover, when examining the marginalized posterior distributions, the influence of the prior can shift the location of the posterior peak away from the true injected values.. These effects can be quantified as a “distance” between the median of the 1-D marginalized distribution and the injected parameter, that is, the one characterizing the simulated source. However, this distance should provide statistical information, since we want to know if the true parameter is contained within a certain credible interval. This also depends on the measurement precision, which is reflected in the spread of the distribution. Given a parameter $\theta^{i}$, a good statistical distance can be defined as the absolute difference between the quantile order related to the injected value and $0.5$, that is, the quantile order of the median value. Given a simulation with injected parameter $\theta^{i}_{\text{inj}}$ and estimated median $\theta^{i}_{\text{med}}$, we thus define the following statistical distance:
@@
where $p(\theta^{i}|s)$ is the posterior marginalized over all the parameters except $\theta_{i}$, and $P(\theta^{i}|s)$ is the cumulative of $p(\theta^{i}|s)$. This quantity approaches zero as the accuracy of the estimate increases. This quantity approaches zero as the estimated parameter gets closer to the injected one. Conversely, bad recoveries would push this quantity toward $0.5$. Whenever an injected parameter is at the boundary of the typical prior support for that variable, it does not make sense to use this distance, since it will always return $0.5$. For instance, when injecting with an inclination angle of $0\degree$, a more indicative quantity might be the, e.g., the 90% credible upper limit (or the lower limit, depending on the cases).
@@ IV.2 Role of the alignment factor
$\mathcal{M}$ Uniform in $0.95\mathcal{M}_{\text{inj}},1.05\mathcal{M}_{\text{inj}}]$
$q$ Uniform in [0.5, 1][0.5,\,1]
@@
We choose slightly different component masses for computational reasons, and we sample over (detector-frame) chirp mass and mass ratio, as these are estimated more directly from GW data. Chirp mass, in particular, is the best-estimated parameter and therefore, to make the simulations faster, we choose a uniform prior centered in its injected value and with a 10% width around it. The injected mass ratio is close to unity, and the prior is taken to be uniform between $0.5$ and $1$. We provide priors for the (dimensionless) spin components aligned with the orbital angular momentum according to Ref. [73]. The extrinsic parameters listed in Table 1 are unchanged throughout these simulations.
The injected inclination angle $\iota$ varies over the simulations: it is taken to give face-on, edge-on, or intermediate-inclination binaries. The prior associated to this quantity is uniform in the cosine of $\iota$. Finally, we use a uniform-in-volume prior for the luminosity distance, which is degenerate with the inclination angle; the support for this prior abundantly contains the injected distance, but it necessarily varies with the injected $d_{L}$ as this is modified to yield different values of optimal SNR.
@@ IV.2.1 Two Advanced LIGO detectors
We now compare these results to the ones for $\hat{\rho}=40$ (bottom row in Fig. 4). One can immediately check that the poor recovery for the two systems at $\hat{\rho}=12$ had different causes. The face-on binary on the left maintains a degenerate profile which is smoother this time around, with a small enhancement in the precision [^8] We can quantify the precision of the estimate as the ratio between $\Delta d_{L}$, that is the larger part of the 90% credible interval, given the asymmetry of the distributions, and the median value $d_{L}^{\text{med}}$. We call it $\Delta d_{L}/d_{L}$ for simplicity. $\Delta d_{L}/d_{L}$ from $36\%$ to $26\%$, and with the upper limit on the inclination angle dropping from $70\degree$ to $55\degree$. The right plot shows the edge-on case, and proves that $\hat{\rho}=40$ is enough to properly appreciate the difference between the two polarizations and to improve the recovery: the injected inclination is almost coincident with the median value, and is included in a 90% credible interval $84\degree,96\degree]$. As far as the distance is concerned, we find $\delta_{\text{st}}(d_{L})=0.26$, so that the injection is inside the 68% credible interval, with a moderate precision of $\Delta d_{L}/d_{L}=16\%$, which would further improve with a larger SNR.
@@ IV.2.3 The importance of signal-to-noise ratio and alignment factor
The other key factor in breaking the degeneracy is represented by the alignment factor $\alpha$, the impact of which has been widely discussed. We perform three simulations of the same system, with and inclination of 66°and detected at an optimal SNR $\hat{\rho}=18.5$ by three different networks: we add ET to the two networks already used so far, but at a different location with respect the one shown in Table 1, chosen in such a way to obtain a frequency-averaged $\alpha\simeq 0.9$. This allows us to have a complete range from $0.1$ to $0.9$. Figure 7 reports the contours of the 68% and 90% credible regions. The injection is inside the 90% credible region in the low-$\bar{\alpha}$ case and enters the 68% region when $\bar{\alpha}$ increases. The main differences between the three configurations can be observed in Table 3. The precision of the estimates, shown in the third and sixth columns, improves for the inclination and seems to worsen for the distance, but this is not an issue, since we worked at fixed SNR. [^9] This implies that the values of the distance were different such to have the same SNR. On the other hand, the goodness of the recovery, described by the other columns, improves significantly, with the difference between the injected inclination and its median estimated value that drops from 48% to 11%. The values are slightly higher for the distance as discussed previously, but the enhancements are significant: from 65% with Advanced LIGO HL to 26% with ET. We have thus proved that the configurations in which the inclination angle is quite far from zero are manageable with a combination of $\bar{\alpha}$ and $\hat{\rho}$ achievable by the upcoming observing runs. However, the recovery of less inclined systems remains a problem difficult to solve even for 3G detectors, unless the SNRs reach very high and rare values [50], or minor effects contribute to distinguishing the two polarization amplitudes and breaking the correlation.
@@ IV.3.2 Low-mass-ratio binaries: an optimistic configuration
We now analyze the case of $\hat{\rho}=40$, when the SNR is twice the one seen so far, using IMRPhenomXHM for the recovery. The 2D marginalized posteriors for $d_{L}$ and $\iota$ are shown in Fig. 9. As expected, the credible regions shrink by a factor $\sim(40/20)^{2}=4$. The difference between injected and recovered distance is about $0.4\%$ for the higher SNR scenario, while the upper limit on the inclination angle drops to $\sim 10\degree$ in this same case. The statistical distance $\delta_{\text{st}}(d_{L})$ remains almost unchanged, while the precision is more than doubled. We summarize the results of these simulations in Table 4.
@@ IV.3.3 The role of ET design for low-mass-ratio binaries
[…]
@@
Figure 10: Joint $d_{L}$$\iota$ posterior distributions for different ET designs, with injected $\iota=0\degree$ (left panels) and $\iota=66\degree$ (right panels); the recovery in the top and bottom panels is performed without and with higher-order modes, respectively. The black lines mark the injected values. The injected luminosity distance is fixed to 4 Gpc, while the SNR values are reported in Table 5.
For each design, we set the inclination angle of the simulated source to two different values: $\iota\in\{0\degree,66\degree\}$. In both cases, we perform a recovery with and without high-order modes and the distance and sky location are kept unchanged. The corresponding optimal SNR values are reported in Table 5.
@@ IV.3.4 Next-to-unity mass ratio binaries I: a more challenging configuration
We begin by taking face-on BBHs with injected (detector-frame) masses of $36\,M_{\odot}$ and $29\,M_{\odot}$, close to the GW150914 values; we also perform a simulation at the same mass ratio but with neutron-star-compatible masses in order to confirm the similarity of the outcomes. For the BBHs, we varied the SNR from 20 to 150. As expected, the impact of high-order modes on the recovery is reduced, given that they are suppressed in the injected signal by the quasi-symmetry of the masses and therefore higher SNRs are appreciate them. The trend of the 90% credible upper limit on the inclination angle as a function of the optimal SNR is reported in the upper panel of Fig. 11, while the lower panel is the posterior distribution for the specific simulation with an optimal SNR of 100. The upper limit value on the inclination angle for $\hat{\rho}=150$ is almost the same as the one estimated at an SNR of 20 but with $q_{\rm inj}=0.1$ (see Table 4). The percentage difference between injected distance and recovered median value passes from 17% to 3%. However, the posterior distribution over the distance, with increasing SNR, does not tend to a normal distribution, but it shrinks asymmetrically toward the injected value that remains out of the 68% credible interval, as shown in the lower panel of Fig. 11. For this reason, the statistical distance $\delta_{\text{st}}(d_{L})$ does not improve.
Figure 12 reports the difference induced by the usage of IMRPhenomXHM or IMRPhenomXP for the recovery, to emphasize that higher-order modes remain important, but their impact is limited by the high mass ratio. The 90% contours are related to an SNR $\hat{\rho}=100$, and the key differences are summarized in Table 6. This table shows that considering high-order modes remains beneficial for both the goodness of the recovery and the precision, but the comparison with Table 4 highlights that a lower mass ratio is crucial.
Moving to BNS systems, we considered masses of $1.4\,M_{\odot}$ and $1.1\,M_{\odot}$, maintaining the same optimal SNR $\hat{\rho}=100$, that for ET corresponds to an injected luminosity distance of $\sim 200$ Mpc. We again used the two different approximants for the recoveries in two separate runs to ensure consistency with the previous test. We do not report the marginalized posteriors, which are analogous to the ones in Fig. 11, but we rather present Table 6 to be compared to Table 7.